OpenStudy (anonymous):

I have to do a check to make sure this is right. When I turned it in she told me my check was wrong. Please show me how to check my answer. Then I have to substitute my answer into the original equation.3b-2=7b+4 This is how I done my problem. Subtract 3b from both sides -2=4b + 4 Subtract 4 from both sides -6=4b Divide both sides by 2 -6 divided by 2= -3 4 divided by 2 = 2 Answer: b=-3/2

OpenStudy (anonymous):

ur solution is exaclty rite :)

OpenStudy (anonymous):

just plug -3/2 everywhere use see b in the equation

OpenStudy (anonymous):

but how do I check it. She wants me to write it out.

OpenStudy (anonymous):

check it like this: lhs:3(-3/2)-2=-13/2 rhs:7(-3/2)+4=-13/2 so both sides are equal and hence verified.

OpenStudy (anonymous):

got it becca ?